Five hundred feet of fencing is available to enclose a rectangular lot along side of highway 65. Cal Trans will supply the fencing for the side along the highway, so only three sides are needed as seen below. What dimensions will produce an area of 40,000 ft2? What is the maximum area that can be enclosed?
Answer by checkley77(12844)

500/4=125 ft. would be the two short sides.
2*125=250 ft. would be the length.
125+250=31,250 ft^2 is the max. area.
Proof:
124*252=31.248 ft^2.
126*248=31,248 ft^2.
A 40,000 ft^2 area is not possible with only 500 ft. of fencing for the three sides of a rectangle lot.